OpenStudy (anonymous):

Angle R = 49º and angle S = 41º. The measure of angle S is greater than the measure of angle R. Angle R and angle S are right angles. Angle R =45º and angle S = 45º.

OpenStudy (anonymous):

7x+3+2x+13=90 solve for x and see what you get for <r and < s

OpenStudy (anonymous):

7x+3x+2x+13=90

OpenStudy (anonymous):

I get 8.22222222. I dont think thats right. 7x nd 2x are like terms = 9x nd so are 3 nd 13 = 16. 9x + 16 = 90 and i get 8.2222.

OpenStudy (anonymous):

are you sure your question says <R=7x+3x and <S= 2x+13

OpenStudy (anonymous):

Yeahh

OpenStudy (anonymous):

then use 7x+3x+2x+13=90, solve for x now

OpenStudy (anonymous):

I get 7

OpenStudy (anonymous):

well i am getting something different 12x=77 x=6.41 isn't it?

OpenStudy (anonymous):

put x= 6.41 and check out your answer

OpenStudy (anonymous):

But dont you combine 7x and 3x which is = 9x and then 2x = 11x.. 11x + 13 = 90. 90 - 13 = 77 / 11 = 7.. alright i'll try that (:

OpenStudy (anonymous):

see your question is <R= 7x+3x and <S= 2x+13 right?

OpenStudy (anonymous):

Yea

OpenStudy (anonymous):

and <R and <S are complementary ..right?

OpenStudy (anonymous):

Yes cuz it = 90

OpenStudy (anonymous):

now 7x+3x+2x+13 should be equal to 90 right 7x+3x+2x+13=90..okay?

OpenStudy (anonymous):

12x+13=90 ..now solve for X lil

OpenStudy (anonymous):

@LilVeroDez13 you have made a mistake in combining 7x and 3x, it's 10x not 9x :)

OpenStudy (anonymous):

Ahh i see! Ur right.. whoops ;) guess i dont know how to add haha ;D

OpenStudy (anonymous):

no it's okay..now solve for x and see are you getting the same answer?

OpenStudy (anonymous):

I get 6.41.. wat u got

OpenStudy (anonymous):

the same

OpenStudy (anonymous):

now find <R and <S

OpenStudy (anonymous):

none of your option seems correct to me

OpenStudy (anonymous):

I know this problem is werid.

OpenStudy (anonymous):

as per the value of x we get <R= 64.1 and <S=90-64.1=25.9